OpenStudy (anonymous):

Find the value of X. Round to the nearest hundredth. 9^(2x+3) = 568

OpenStudy (anonymous):

real solution -0.06

OpenStudy (anonymous):

x=-0.056788076162439606408046041536045944

OpenStudy (anonymous):

Thanks, but i'm stuck in the work, i'm not sure how one would get there. Could you show the work for me please ;)?

OpenStudy (anonymous):

Ok you have \[9^{(2x+3)}=568\] take the log of both sides to base 9, so you have \[2x+3 = \log_9(568)\rightarrow \frac{\ln{568}}{\ln(9)}\] subtract 3 from both sides to get \[2x = \frac{\ln{568}-3\ln9}{\ln(9)} \Rightarrow x= \frac{\ln{568}-3\ln9}{2\ln(9)}\] which is approximately the real solution \[x =-0.056788076162439606408046041536045\]

OpenStudy (anonymous):

\[\ln x\] is the natural log of x